Operations Report – May 19th

Crew 299 Operations Report 5-19-2024

SOL: 7

Name of person filing report: Rishabh Pandey (Engineer)

Non-Nominal Systems: Suit 9 (broken visor), Suit 11 (fan issue)

Notes on non-nominal systems: Issues with both suits were noted before our mission, both will not be used during our mission.

Rovers Used:
            Hours: 282.2
            Beginning charge: 100%
            Ending charge: 50%
            Currently charging: yes
            Hours: 189.8
            Beginning charge: 100%
            Ending charge: 53%
            Currently charging: yes
General notes on rovers: None

Summary of Hab operations:
      WATER USE: 50 Gallons
      Water (static tank): 192.3 (using effective value from spreadsheet)
      Static tank pipe heater: off
      Static tank heater: off
      Toilet tank emptied: no

Summary of internet usage: Catch up on Earthly news, email responses, code development, entertainment

Summary of suits and radios: Used Radios: 1,2,3,6, Suits: 6,7,8,10

Summary of GreenHab operations: NA, GreenHab out of operation

Summary of ScienceDome operations: More rock samples were made and analyzed. Radishes were watered and checked in on.

Summary of RAM operations: Multimeter and Measuring tape used
Summary of observatory issues: NA

Copyright © The Mars Society. All rights reserved. | Main Site